Mr Gorman spent $18 on some paintbrushes and colored pencils. He bought 4 times as many colored pencils as paintbrushes. Each paintbrush cost $3 and and each colored pencil cost $0.75. How many colored pencils did Mr. Gorman buy?
Found 3 solutions by lynnlo, ikleyn, timofer:
Answer by lynnlo(4176)
Answer by ikleyn(53421)

.
Mr. Gorman spent $18 on some paintbrushes and colored pencils.
He bought 4 times as many colored pencils as paintbrushes.
Each paintbrush cost $3 and each colored pencil cost $0.75.
How many colored pencils did Mr. Gorman buy?
~~~~~~~~~~~~~~~~~~~~~


        This problem ideally suits to be solved mentally.


In accordance with the problem, we can group colored pencils and paintbrushes in sets,
each containing 4 colored pencils and one paintbrush.


Every such group costs 4*0.75 + 3 = 6 dollars, giving the total of 18 dollars.


Hence, the number of such groups is  18/6 = 3.


Hence, the number of colored pencils was  4*3 = 12.    ANSWER

Solved.

You can do algebra in the same way - it will lead you to the same answer.



Answer by timofer(136)

This problem is too much like the fundraiser problem. Solve this in the same way.